# Architectural Co-Design for Zero-Shot Anomaly Detection: Decoupling Representation and Dynamically Fusing Features in CLIP

### v2

@@ Abstract
Pre-trained Vision-Language Models (VLMs) face a significant adaptation gap when applied to Zero-Shot Anomaly Detection (ZSAD), stemming from their lack of local inductive biases for dense prediction and their reliance on inflexible feature fusion paradigms. We address these limitations through an Architectural Co-Design framework that jointly refines feature representation and cross-modal fusion. Our method integrates a parameter-efficient Convolutional Low-Rank Adaptation (Conv-LoRA) adapter to inject local inductive biases for fine-grained representation, and introduces a Dynamic Fusion Gateway (DFG) that leverages visual context to adaptively modulate text prompts, enabling a powerful bidirectional fusion. Extensive experiments on diverse industrial and medical benchmarks demonstrate superior accuracy and robustness, validating that this synergistic co-design is critical for robustly adapting foundation models to dense perception tasks.
@@
Fig. 1: Comparison of fusion paradigms. (a) Prior works rely on a rigid, static alignment between corresponding feature blocks. (b) Our Architectural Co-Design enables a flexible fusion policy by first enriching visual features with local priors (Conv-LoRA) and then dynamically generating tailored text descriptors for each visual level (DFG).
Fig. 2: Overview of the proposed ACD-CLIP architecture. (a) The Overall Framework: We structure CLIP’s vision and text encoders into a hierarchy NN sequential Groups (as illustrated, $N=3$). Each vision group is enhanced by a trainable Conv-LoRA Adapter to instill local priors. The Dynamic Fusion Gateway (DFG) then uses each visual feature $V_{i}$ to generate a tailored text descriptor for producing a level-specific anomaly map $M_{i}$. (b) The Conv-LoRA Adapter: Our parameter-efficient adapter features a multi-branch design with multi-scale convolutions inside a LoRA bottleneck.
@@ 1 Introduction
Zero-Shot Anomaly Detection (ZSAD) leverages Vision-Language Models (VLMs) [1, 2] like CLIP [3] to bypass the need for extensive, category-specific training data required by traditional methods [4, 5]. The dominant paradigm adapts these VLMs via text prompts, which has evolved from hand-crafted ensembles (e.g., WinCLIP [6]) to learnable prompts tuned on auxiliary data. These advanced methods explore object-agnostic semantics (AnomalyCLIP [7]), dynamic prompt generation (AdaCLIP [8]), and multi-layer feature queries (CLIP-AD [9]). While significantly advancing prompt engineering, this line of work faces two fundamental limitations.
@@
To address these core issues, we introduce the Architectural Co-Design CLIP (ACD-CLIP) framework [Figure 1(b)], a novel approach that synergistically refines both feature representation and cross-modal fusion. The synergy is crucial: our adapter provides the fine-grained visual details that static fusion would overlook, while the gateway supplies the adaptive mechanism required to fully exploit them. Our architectural modifications are enabled by Parameter-Efficient Fine-Tuning (PEFT) [13, 14]. While recent works have shown the value of integrating convolutional structures into PEFT adapters for general vision tasks [15, 16], our work is the first to co-design such an adapter with a dynamic fusion mechanism specifically for ZSAD. Our contributions are: 1) A parameter-efficient Conv-LoRA Adapter to inject local inductive biases for fine-grained representation. 2) A Dynamic Fusion Gateway, a vision-guided mechanism enabling a flexible, bidirectional fusion policy. 3) Validation of our synergistic co-design, which yields significant performance gains on diverse ZSAD benchmarks.
@@ 2 Methodology
As illustrated in Figure 2, our ACD-CLIP framework introduces two core innovations to adapt VLMs for ZSAD: (1) an architectural co-design for decoupled representation learning, and (2) a dynamic gateway for cross-modal fusion. We detail these components below.
@@ 2.1 Hierarchical Feature Adaptation with Local Priors
The adapter features a parallel multi-branch design, where each branch employs a distinct convolutional kernel size $k\times k$ for $k\in\{3,5\}$, operating within a LoRA bottleneck. This allows it to capture local patterns at varied receptive fields. The entire operation of the adapter, denoted as $\Phi_{\text{Adapter}}$, is formulated to generate a residual update $\Delta X$ as follows:
@@
The output of each branch, $X_{\text{branch\_out}}^{(k)}$, is generated in a two-step process. First, the input tensor $X_{\text{in}}$ is passed through the LoRA down-projection and a strided down-convolution to produce an intermediate downsampled feature map, $X_{\text{down}}^{(k)}$:
$X_{\text{down}}^{(k)}=\operatorname{Conv}_{\text{down}}^{(k)}\left(\mathcal{R}_{\text{2D}}(X_{\text{in}}W_{\text{down}})\right)$ (2)
This feature map is then upsampled via a transposed convolution, reshaped, and passed through the LoRA up-projection matrix to yield the final branch output:
$X_{\text{branch\_out}}^{(k)}=\mathcal{R}_{\text{seq}}\left(\frac{1}{k^{2}}\operatorname{Conv}_{\text{up}}^{(k)}\left(X_{\text{down}}^{(k)}\right)\right)W_{\text{up}}$ (3)
Here, $W_{\text{down}}$ and $W_{\text{up}}$ are the LoRA projection matrices. $\mathcal{R}_{\text{2D}}$ and $\mathcal{R}_{\text{seq}}$ are operators for reshaping between sequence and spatial formats. $\operatorname{Conv}_{\text{down}}^{(k)}$ is a strided convolution and $\operatorname{Conv}_{\text{up}}^{(k)}$ is a transposed convolution.
@@ 3.2 Main Results and Analysis
As shown in Table 1, our ACD-CLIP framework consistently establishes a new state-of-the-art across both industrial and medical domains. On the widely-used MVTec-AD benchmark, our model demonstrates a clear advantage in fine-grained localization, outperforming methods like AnomalyCLIP in pixel-level AP by a significant margin of nearly +10%. This superiority in capturing precise anomaly boundaries is a direct result of our architectural co-design.
Crucially, our approach shows strong cross-domain generalization. When trained solely on industrial data, ACD-CLIP excels on challenging medical benchmarks, achieving a pixel-level AUROC of 90.4% on ClinicDB and 96.6% on BrainMRI. These quantitative gains are supported by qualitative results (Figure 3), which show our model generates significantly cleaner and more accurate anomaly maps with less background noise compared to other baselines.
@@ 3.3 Ablation Study
Our ablation study (Table 2) validates the individual and synergistic contributions of our two core components against a baseline using a linear adapter and static fusion. Integrating the Conv-LoRA Adapter alone boosts the pixel-level AUROC by 6.8%, confirming its critical role in providing the local inductive biases necessary for dense prediction. Separately, adding the Dynamic Fusion Gateway (DFG) improves the image-level AUROC by 4.7%, demonstrating the effectiveness of its vision-guided, adaptive fusion mechanism. The full ACD-CLIP model, which combines both, achieves the highest performance, confirming a powerful synergistic effect between enhancing feature representation and enabling dynamic cross-modal fusion.

### v3

@@ Abstract
Pre-trained Vision-Language Models (VLMs) face a significant adaptation gap when applied to Zero-Shot Anomaly Detection (ZSAD), stemming from their lack of local inductive biases for dense prediction and their reliance on inflexible feature fusion paradigms. We address these limitations through an Architectural Co-Design framework that jointly refines feature representation and cross-modal fusion. Our method proposes a parameter-efficient Convolutional Low-Rank Adaptation (Conv-LoRA) adapter to inject local inductive biases for fine-grained representation, and introduces a Dynamic Fusion Gateway (DFG) that leverages visual context to adaptively modulate text prompts, enabling a powerful bidirectional fusion. Extensive experiments on diverse industrial and medical benchmarks demonstrate superior accuracy and robustness, validating that this synergistic co-design is critical for robustly adapting foundation models to dense perception tasks.
@@
Fig. 1: Comparison of fusion paradigms. (a) Prior works rely on a rigid, static alignment between corresponding feature blocks. (b) Our Architectural Co-Design enables a flexible fusion policy by enriching visual features with local priors (Conv-LoRA) and then dynamically generating tailored text descriptors for each visual level (DFG).
Fig. 2: Overview of the proposed ACD-CLIP architecture. (a) The Overall Framework: We structure CLIP’s vision and text encoders into a hierarchy of NN sequential Groups (as illustrated, $N=3$). Each vision group is enhanced by a trainable Conv-LoRA Adapter to instill local priors, while each corresponding text group incorporates a standard LoRA adapter. The Dynamic Fusion Gateway (DFG) then uses each visual feature $V_{i}$ to generate a tailored text descriptor for producing a level-specific anomaly map. (b) The Conv-LoRA Adapter: Our parameter-efficient adapter features a multi-branch design with multi-scale convolutions inside a LoRA bottleneck.
@@ 1 Introduction
Zero-Shot Anomaly Detection (ZSAD) leverages Vision-Language Models (VLMs) [1, 2] like CLIP [3] to bypass the need for extensive, category-specific training data required by traditional methods [4, 5]. The dominant paradigm adapts these VLMs via text prompts, which has evolved from hand-crafted ensembles (e.g., WinCLIP [6]) to learnable prompts tuned on auxiliary data. These advanced methods explore object-agnostic semantics (AnomalyCLIP [7]), dynamic prompt generation (AdaCLIP [8]), and multi-layer feature queries (CLIP-AD [9]). However, this line of work faces two fundamental limitations.
@@
To address these core issues, we introduce the Architectural Co-Design CLIP (ACD-CLIP) framework [Figure 1(b)], a novel approach that synergistically refines both feature representation and cross-modal fusion. The synergy is crucial: the Conv-LoRA adapter provides fine-grained details that are fully exploited by the DFG’s adaptive fusion mechanism. Our architectural modifications are enabled by Parameter-Efficient Fine-Tuning (PEFT) [13, 14]. While recent works have shown the value of integrating convolutional structures into PEFT adapters for general vision tasks [15, 16], our work is the first to co-design such an adapter with a dynamic fusion mechanism specifically for ZSAD. Our contributions are: 1) A parameter-efficient Conv-LoRA Adapter to inject local inductive biases for fine-grained representation. 2) A Dynamic Fusion Gateway, a vision-guided mechanism enabling a flexible, bidirectional fusion policy. 3) Validation of our method, which yields significant performance gains on diverse industrial and medical ZSAD benchmarks.
@@ 2 Methodology
As illustrated in Figure 2, our ACD-CLIP framework introduces two core innovations to adapt VLMs for ZSAD: (1) an architectural co-design for decoupled representation learning, and (2) a dynamic gateway for cross-modal fusion.
@@ 2.1 Hierarchical Feature Adaptation with Local Priors
The adapter features a parallel multi-branch design with distinct $k\times k$ convolutional kernels (where $k\in\{3,5\}$) that operate within a LoRA-style bottleneck. This structure allows it to capture local patterns at varied receptive fields while maintaining parameter efficiency. The entire operation, $\Phi_{\text{Adapter}}$, generates a residual update $\Delta X$ by fusing the outputs of the parallel branches:
@@
The output of each branch, $X_{\text{branch\_out}}^{(k)}$, is generated in a two-step process. First, the input tensor $X_{\text{in}}$ is projected to a low-rank space by $W_{\text{down}}$, reshaped, and passed through a $k\times k$ convolution to instill local context and transform the features, producing an intermediate feature map $X_{\text{conv}}^{(k)}$:
$X_{\text{conv}}^{(k)}=\frac{1}{k}\operatorname{Conv}_{\text{down}}^{(k)}\left(\mathcal{R}_{\text{2D}}(X_{\text{in}}W_{\text{down}})\right)$ (2)
This feature map is then processed by a second $k\times k$ convolution for further feature refinement, reshaped back into a sequence, and projected up by $W_{\text{up}}$ to yield the final branch output:
$X_{\text{branch\_out}}^{(k)}=\mathcal{R}_{\text{seq}}\left(\frac{1}{k}\operatorname{Conv}_{\text{up}}^{(k)}\left(X_{\text{conv}}^{(k)}\right)\right)W_{\text{up}}$ (3)
Here, $W_{\text{down}}$ and $W_{\text{up}}$ are the LoRA projection matrices performing channel-wise compression and expansion. $\mathcal{R}_{\text{2D}}$ and $\mathcal{R}_{\text{seq}}$ are operators for reshaping between sequence and spatial formats. $\operatorname{Conv}_{\text{up}}^{(k)}$ and $\operatorname{Conv}_{\text{down}}^{(k)}$ are standard $k\times k$ convolutions that operate on feature channels for transformation and local context infusion while preserving spatial dimensions.
@@ 3.2 Main Results and Analysis
As shown in Table 1, our ACD-CLIP framework consistently establishes a new state-of-the-art across both industrial and medical domains. On the widely-used MVTec-AD benchmark, our model excels at fine-grained localization, boosting pixel-level AP by nearly +10% over AnomalyCLIP. This superiority in capturing precise anomaly boundaries is a direct result of our architectural co-design.
Crucially, our approach shows strong cross-domain generalization. When trained solely on industrial data, ACD-CLIP excels on challenging medical benchmarks, achieving a pixel-level AUROC of 90.4% on ClinicDB and 96.6% on BrainMRI. Qualitative results (Figure 3) corroborate these gains, showing our model produces cleaner, more accurate anomaly maps with less noise.
Interestingly, our results across different numbers of groups ($N$) show that performance generally peaks around $N=3$, suggesting a trade-off between hierarchical specialization and model complexity, where deeper partitioning does not necessarily yield further gains and may lead to overfitting on the auxiliary dataset.
@@ 3.3 Ablation Study
Our ablation study (Table 2) validates the individual and synergistic contributions of our two core components against a baseline using a standard LoRA adapter on ViT and static fusion. Integrating the Conv-LoRA Adapter alone boosts the pixel-level AUROC by 6.8%, confirming its critical role in providing the local inductive biases necessary for dense prediction. Separately, adding the Dynamic Fusion Gateway (DFG) improves the image-level AUROC by 4.7%, demonstrating the effectiveness of its vision-guided, adaptive fusion mechanism. The full ACD-CLIP model, which combines both, achieves the highest performance, confirming a powerful synergistic effect between enhancing feature representation and enabling dynamic cross-modal fusion.